# SMARTe: Slot-based Method for Accountable Relational Triple extraction

### v1

@@ 1 Introduction
Relational Triple Extraction (RTE) is a well-established and widely studied task in Natural Language Processing (NLP). Its primary objective is to automatically extract structured information such as names, dates, and relationships from unstructured text, thereby enhancing data organization and accessibility Nayak et al. (2021). These extracted relationships are represented as relational triples consisting of (Subject, Relation, Object). For example, the sentence “Barack Obama is the President of America” can be expressed as the triple (Barack Obama, Head of, America). Such structured representations underpin a range of downstream applications, including knowledge graph construction, question answering, and information retrieval. RTE is often confused with Joint Entity and Relation Extraction (JERE) Zhang et al. (2017); Gupta et al. (2016); Miwa and Sasaki (2014), or the terms are used interchangeably by some scholars Sui et al. (2023); Li et al. (2021). However, there is a key distinction: JERE addresses both entity identification and relation extraction, which require annotation of all entities in a text, regardless whether they participate in a relationship. In contrast, RTE focuses exclusively on entities that are part of a relationship, specifically a set of annotated relational triples. In this paper, we focus solely on RTE.
Traditional approaches to RTE often prioritize performance metrics such as precision, recall, and F1 score, with limited consideration for interpretability. According to a comprehensive survey by Zhao et al. (2024), we found that many existing methods introduce novel architectures, claiming that these designs facilitate specific interactions that enhance model performance. However, such claims are frequently supported solely by empirical results, with minimal analysis or explanation of the underlying mechanisms, leaving the validity of these assertions largely unexplored. Moreover, we observed that none of the reviewed methods explicitly address explainability, which highlights a significant gap in current research. In this study, we introduce SMARTe111Our implementation can be found at https://github.com/Chen-XueWen/SMARTe: a Slot-based Method for Accountable Relational Triple Extraction, a transparent architecture designed to address interpretability in RTE. SMARTe incorporates intrinsic interpretability through the use of a slot attention mechanism. This approach ensures that every prediction from SMARTe can be explicitly traced back to its learned slot representations and the specific tokens contributing to each predicted relational triple. This results in informed and transparent outputs, embodying the Accountability, which directly equates to explainability, in our model. Since the learned slots do not follow any inherent order, the RTE task is framed as a set prediction problem.
@@
Our goal is to demonstrate how slot attention can be effectively applied to NLP tasks to generate meaningful explanations. By clustering relevant information, this method facilitates informed and transparent predictions. Although we illustrate our approach in the context of RTE, it is broadly applicable to other NLP tasks characterized by set-based target structures. Interpretability becomes especially valuable in high-stakes scenarios where transparency and trust are paramount Danilevsky et al. (2020). As stakeholders increasingly demand clear and interpretable explanations for AI-driven decisions, our approach effectively addresses these critical requirements. We introduce SMARTe, a Slot-based Method for Accountable Relational Triple extraction. Our contributions are as follows:
@@
To the best of our knowledge, the proposed SMARTe framework is the first to introduce interpretability to relational triple extraction tasks, marking a significant contribution to the field of Explainable AI (XAI) in Natural Language Processing (NLP). This is achieved by adapting the slot attention mechanism, originally developed for unsupervised learning in computer vision.
@@ 2 Related Work
Our approach builds on the seq2seq paradigm, eliminating the need to permute token pairs. Slot attention also disentangles relational triplets into distinct slots, delivering superior performance over the earlier seq2seq and tagging-based models. It is also competitive to pairwise-based methods while maintaining computational efficiency. Most importantly, the RTE field has given limited attention to explainability. Existing research primarily focuses on optimizing performance metrics such as F1 scores and surpassing benchmarks, often achieving only marginal gains (typically less than 1%, as seen in newer research). Since benchmark datasets already exhibit performance exceeding 90%, this raises questions about their practical relevance. The latest SOTA models (usually based on pairwise methods) lack interpretability, which stems from their prediction mechanisms that classify relationships by evaluating token pairs in a pairwise manner. Since these models rely on token representations derived from fine-tuned BERT models, the surrounding context is already embedded within these token pairs, making it difficult to trace the specific context each pair has “absorbed.”. For example, given the sentence “Barack Obama is the President of America,” these models predict the triple (Barack Obama, Head of, America) by relying solely on the representations of Barack Obama and America to classify the relationship Head of. The token president does not contribute to the prediction, as its information is absorbed into the representations of the paired tokens. In contrast, our model employs slot attention to directly learn relational triple representations. This method allows us to explicitly identify the tokens contributing to each slot, which corresponds to a relational triple prediction. By focusing on slot-level representations rather than pairwise token interactions, our approach enhances interpretability, as it provides a clearer understanding of how individual tokens influence the final predictions from the slot attention mechanism as shown in Appendix A.
@@ 4 Experimental Setup
We evaluate our model using two widely recognized benchmark datasets frequently employed in the existing literature on relational triple extraction: the New York Times (NYT) dataset Riedel et al. (2010) and the Web Natural Language Generation (WebNLG) dataset Gardent et al. (2017). Both datasets include two variants of annotation: Partial Matching and Exact Matching. In Partial Matching, only the head words of the ground truth entities are annotated, whereas in Exact Matching, the entire span of the entities is annotated. Therefore, Exact Matching offers a more precise and comprehensive representation of the task. It is worth mentioning that most prior studies primarily report results based on Partial Matching, often neglecting Exact Matching. In our study, we present results for both matching strategies to facilitate comparisons in future research. Detailed statistics for these datasets are provided in Table 5 and 6. Please note that these are the only two benchmark datasets specifically designed for this task. Our experiments were conducted on a GeForce RTX 2080 Ti GPU (11GB VRAM) and ran with 12 random seeds (11, 29, 33, 39, 42, 53, 57, 62, 65, 73, 96, 98). Training time per epoch is approximately 7 minutes for NYT and 1 minute for WebNLG. Please refer to Appendix C for a detailed overview of the key hyperparameters used in our experiments.
@@ 5 Experimental Results
Our results are benchmarked against the latest state-of-the-art (SOTA) methods from 2020 onwards, as we find comparisons with outdated models serve limited purpose beyond formality. To ensure a fair comparison, all models including ours, use BERT-Base-Cased as encoder, with all comparative results sourced directly from their original publications. For our model, we report the mean and standard deviation across multiple runs with our provided seeds in our previous section. Moreover, we also include the best-performing seed in terms of overall F1 score for reference (indicated with ), recognizing that many related works do not explicitly state whether their results are derived from experiments using multiple random seeds or from optimizing the F1 score with a single specific seed. Furthermore, we include results for the softmax variant as a comparison to the optimal transport approach (labeled as “Opt Transport”). Our SMARTe model achieves highly competitive performance, coming within 1% of the top-performing methods and outperforming most previous attempts, while offering the added advantage of interpretability. We also acknowledge that our model may exhibit a slightly higher dependency on larger datasets compared to SOTA methods as this became evident in the WebNLG exact matching benchmark, where our model’s performance was comparatively lower. We attribute this to the presence of numerous poorly defined relationships in the dataset, as illustrated in Figure 11, many of which have fewer than 10 training examples. Additionally, it is also important to note that many previous models were not benchmarked against this variant, making the comparison less conclusive. Furthermore, our analysis shows that the optimal transport algorithm consistently outperforms the softmax approach, particularly in handling overlapping patterns, where the softmax approach shows significantly weaker performance. These findings justify the integration of optimal transport into our model for enhanced performance.
@@ 6 Qualitative Analysis of Explanations
For location-related relationships, such as /location/location/contains, the attention mechanism predominantly focuses on the object location, as it appears to encapsulate all the necessary information needed to predict the relational triple, as shown in Figure 5 & 6. In these instances, the model’s task shifts from deriving relational meaning from the sentence to recognizing pre-existing factual associations. Since such location-based relationships typically represent fixed world knowledge, the annotations often reflect factual truths rather than information inferred from the sentence context. This underscores a key limitation in providing meaningful explanations for certain types of relationships, as they depend more on external knowledge or the training data than on contextual cues within the text.
@@
By providing clear and interpretable explanations, our SMARTe model establishes a new benchmark in understanding and visualizing relational reasoning, particularly in areas where many existing models fail. Notably, the quality of explanations can vary based on the complexity and nature of the relationship. This is especially evident when dealing with less straightforward or sparsely represented relationships, where limited training data can hinder the model’s ability to produce coherent explanations. Despite these challenges, the model’s capacity to deliver meaningful insights across a wide range of relationships demonstrates its potential for improving interpretability and fostering a deeper understanding of relational extraction tasks.
@@ 7 Conclusion
In this paper, we introduce SMARTe: a Slot-based Method for Accountable Relational Triple extraction, addressing the critical need for interpretability in relational triple extraction models. By leveraging a slot attention mechanism and framing the task as a set prediction problem, SMARTe ensures that predictions are explicitly traceable to their learned representations, providing intrinsic transparency. Our experimental results on the NYT and WebNLG datasets demonstrate that SMARTe achieves performance comparable to state-of-the-art models (within 1% range), while simultaneously offering meaningful explanations through attention heatmaps. These findings underscore the feasibility of combining interpretability with effectiveness, addressing the “black-box” limitations inherent in prior approaches. The pursuit of interpretability in machine learning models is not merely an academic exercise but a pressing necessity. As NLP systems are increasingly deployed in high-stakes domains requiring accountability, such as healthcare Loh et al. (2022), finance Chen et al. (2023), and law Górski and Ramakrishna (2021), ensuring models are both effective and interpretable becomes paramount.
@@ 8 Limitations
For limitations, we acknowledge that our model’s performance slightly lags behind leading models due to its seq2seq design, which does not incorporate combinatorial token interactions like bipartite graphs or interaction tables. Future work will focus on improving token interactions to boost performance while preserving interpretability. Additionally, explanations for complex interactions remain indirect and less intuitive, highlighting differences between model and human text interpretation. To address this, we plan to explore more user-friendly explanation methods to enable users to derive clearer and more actionable insights from predictions.
@@ Appendix C Experiment Hyperparameter Configuration
Appendix D Licensing and Terms of Use
D.1 Dataset
@@
D.2 Model and Code Release
Our SMARTe implementation publicly available at https://github.com/Chen-XueWen/SMARTe.
D.3 Software Dependencies and Implementation
@@
D.4 Ethical and Legal Considerations
@@
D.5 Information About Use Of AI Assistants

### v2

@@ 1 Introduction
Relational Triple Extraction (RTE) is a well-established and widely studied task in Natural Language Processing (NLP). Its primary objective is to automatically extract structured information such as names, dates, and relationships from unstructured text, thereby enhancing data organization and accessibility Nayak et al. (2021). These extracted relationships are represented as relational triples consisting of (Subject, Relation, Object).Such structured representations underpin a range of downstream applications, including knowledge graph construction, question answering, and information retrieval. RTE is often confused with Joint Entity and Relation Extraction (JERE) Zhang et al. (2017); Gupta et al. (2016); Miwa and Sasaki (2014), or the terms are used interchangeably by some scholars Sui et al. (2023); Li et al. (2021). However, there is a key distinction: JERE addresses both entity identification and relation extraction (i.e., ACE04, ACE05 and CoNLL04 datasets) Roth and Yih (2004), which require annotation of all entities in a text, regardless whether they participate in a relationship. In contrast, RTE focuses exclusively on entities that are part of a relationship, specifically a set of annotated relational triples, hence the two tasks are fundamentally different. In this paper, we focus solely on RTE.
To motivate our work, lets take a look at an example seen in Figure 1. Most approaches rely on fine-tuned BERT embeddings to form pair representations by permuting all token pairs and classifying their relationships. However, this methodology makes it difficult to trace which specific contextual information each pair has absorbed, for instance, whether "Barack" or "America" has incorporated the "president" context. Despite achieving superior F1 performance, these pair-based approaches lack interpretability, obscuring how models arrive at their predictions. In contrast, our approach uses triplet-based representation, which absorbs all relevant tokens to directly form relational triplets. This enables a more direct and interpretable encoding of relationships. We observed that most existing methods have excellent performance but with limited interpretability, which highlights a significant gap in current research. In this study, we introduce SMARTe111We will release our implementation after acceptance.: a Slot-based Method for Accountable Relational Triple Extraction, a transparent architecture designed to address interpretability in RTE. SMARTe incorporates intrinsic interpretability through the use of a slot attention mechanism. This approach ensures that every prediction from SMARTe can be explicitly traced back to its learned slot representations and the specific tokens contributing to each predicted relational triple as shown in Appendix A. This results in informed and transparent outputs, embodying the Accountability, which directly equates to explainability, in our model. Since the learned slots do not follow any inherent order, the RTE task is framed as a set prediction problem.
@@
Although we illustrate our approach in the context of RTE, it is broadly applicable to other NLP tasks characterized by set-based target structures. Interpretability becomes especially valuable in high-stakes scenarios where transparency and trust are paramount Danilevsky et al. (2020). As stakeholders increasingly demand clear and interpretable explanations for AI-driven decisions, our approach effectively addresses these critical requirements. We introduce SMARTe, a Slot-based Method for Accountable Relational Triple extraction. Our contributions are as follows:
@@
To the best of our knowledge, the proposed SMARTe framework is the first to introduce interpretability to relational triple extraction tasks by adapting the slot attention mechanism, originally developed for unsupervised learning in computer vision.
@@
We report zero-shot performance on RTE tasks using recent LLMs for future benchmarks.
@@ 4 Experimental Setup
We evaluate our model using two widely recognized benchmark datasets frequently employed in the existing literature on relational triple extraction: the New York Times (NYT) dataset Riedel et al. (2010) and the Web Natural Language Generation (WebNLG) dataset Gardent et al. (2017) (Please note that these are the only two benchmark datasets specifically designed for this task). Both datasets include two variants of annotation: Partial Matching and Exact Matching. In Partial Matching, only the head words of the ground truth entities are annotated, whereas in Exact Matching, the entire span of the entities is annotated. Therefore, Exact Matching offers a more precise and comprehensive representation of the task. It is worth mentioning that most prior studies primarily report results based on Partial Matching, often neglecting Exact Matching. In our study, we present results for both matching strategies to facilitate comparisons in future research. Detailed statistics for these datasets are provided in Table 5 and 6. Please refer to Appendix C for a detailed overview of the key hyperparameters and hardware used in our experiments.
@@ 5 Experimental Results
| Phi4-14B‡ Abdin et al. (2024) | - | - | - | 3.3 | 6.1 | 4.0 |
| Gemma3-27B‡ Gemma-Team (2025) | - | - | - | 6.0 | 12.0 | 7.3 |
| Qwen3-32B‡ Qwen-Team (2025) | - | - | - | 3.7 | 8.2 | 4.7 |
| Llama3.3-70B‡ Grattafiori et al. (2024) | - | - | - | 7.0 | 15.2 | 8.9 |
| GPT 4o-Mini‡ OpenAI et al. (2024) | - | - | - | 3.3 | 6.3 | 3.9 |
| GPT 4.1-Mini‡ OpenAI et al. (2024) | - | - | - | 7.1 | 11.3 | 7.9 |
@@
| Phi4-14B‡ Abdin et al. (2024) | - | - | - | 13.5 | 16.4 | 14.3 |
| Gemma3-27B‡ Gemma-Team (2025) | - | - | - | 24.6 | 31.4 | 26.6 |
| Qwen3-32B‡ Qwen-Team (2025) | - | - | - | 15.7 | 19.2 | 16.8 |
| Llama3.3-70B‡ Grattafiori et al. (2024) | - | - | - | 25.7 | 31.3 | 27.3 |
| GPT 4o-Mini‡ OpenAI et al. (2024) | - | - | - | 24.0 | 28.8 | 25.3 |
| GPT 4.1-Mini‡ OpenAI et al. (2024) | - | - | - | 29.6 | 32.8 | 30.1 |
@@
Our results are benchmarked against the latest state-of-the-art (SOTA) methods from 2020 onwards, as we find comparisons with outdated models serve limited purpose beyond formality. To ensure a fair comparison, all models including ours, use BERT-Base-Cased as encoder, with all comparative results sourced directly from their original publications. For our model, we report the mean and standard deviation across multiple runs with our provided seeds in our previous section. Moreover, we also include the best-performing seed in terms of overall F1 score for reference (indicated with ), recognizing that many related works do not explicitly state whether their results are derived from experiments using multiple random seeds or from optimizing the F1 score with a single specific seed. Furthermore, we include results for the softmax variant as a comparison to the optimal transport approach (labeled as “Opt Transport”). We include zero-shot prompting results from recent LLMs to assess their inherent reasoning capabilities without task-specific fine-tuning, which also reflects a more common and practical usage scenario. Few-shot prompting proves impractical in our context due to extensive relation types (100+ for WebNLG) and multiple relations per text, which complicate prompt design. Fine-tuning LLMs in the same way as our model would offer little advantage over fine-tuning BERT, as it effectively reduces LLMs to encoder backbones and neglects their key strengths such as in-context learning and zero-shot generalization. We therefore include zero-shot LLM results as a reference point rather than a direct competitive comparison, acknowledging the fundamental differences in paradigm and optimization objectives. Note this assessment applies only to Exact matching, as LLMs do not inherently identify the annotated head word required for Partial matching.
Our SMARTe model achieves highly competitive performance, coming within 1% of the top-performing methods and outperforming most previous attempts, while offering the added advantage of interpretability. We also acknowledge that our model may exhibit a slightly higher dependency on larger datasets compared to SOTA methods as this became evident in the WebNLG exact matching benchmark, where our model’s performance was comparatively lower. We attribute this to the presence of numerous poorly defined relationships in the dataset, as illustrated in Figure 12, many of which have fewer than 10 training examples. Additionally, it is also important to note that many previous models were not benchmarked against this variant, making the comparison less conclusive. Furthermore, our analysis shows that the optimal transport algorithm consistently outperforms the softmax approach, particularly in handling overlapping patterns, where the softmax approach shows significantly weaker performance. These findings justify the integration of optimal transport into our model for enhanced performance.
@@ 6 Qualitative Analysis of Explanations
For location-related relationships, such as /location/location/contains, the attention mechanism predominantly focuses on the object location, as it appears to encapsulate all the necessary information needed to predict the relational triple, as shown in Figure 6 & 7. In these instances, the model’s task shifts from deriving relational meaning from the sentence to recognizing pre-existing factual associations. Since such location-based relationships typically represent fixed world knowledge, the annotations often reflect factual truths rather than information inferred from the sentence context. This underscores a key limitation in providing meaningful explanations for certain types of relationships, as they depend more on external knowledge or the training data than on contextual cues within the text.
@@
While explanation quality varies with relationship complexity and characteristics, the model’s ability to provide meaningful insights across diverse relationships demonstrates its potential to enhance interpretability and deepen understanding of relational extraction tasks. SMARTe serves as a valuable stepping stone toward truly interpretable relational reasoning in AI systems. Beyond relational triple extraction, this methodology can also be used for a broad range of set-related NLP tasks.
@@ 7 Conclusion
In this paper, we introduce SMARTe that addresses the critical need for interpretability in relation extraction models. By leveraging slot attention and framing the task as set prediction, SMARTe ensures predictions are explicitly traceable to their learned representations. Our experiments on NYT and WebNLG datasets show that SMARTe achieves performance comparable to state-of-the-art models (within 1% range) while providing meaningful explanations through slots. These findings demonstrate the feasibility of combining interpretability with effectiveness, addressing the "black-box" limitations of prior approaches. This work responds to the growing necessity for interpretable ML models as NLP systems are increasingly deployed in high-stakes domains requiring accountability, including healthcare Loh et al. (2022), finance Chen et al. (2023), and law Górski and Ramakrishna (2021).
@@ 8 Limitations
For limitations, we acknowledge that our model’s performance slightly lags behind leading models due to its seq2seq design, which does not incorporate combinatorial token interactions like bipartite graphs or interaction tables. Future work will focus on improving token interactions to boost performance while preserving interpretability. Additionally, explanations for complex interactions remain indirect and less intuitive, highlighting differences between model and human text interpretation, which is a known issue with attention-based explanation. To address this, we plan to explore more user-friendly explanation methods to enable users to derive clearer and more actionable insights from predictions.
Another limitation is the computational cost of Hungarian matching, which scales at where n represents our budgeted slot count. While this remains manageable for the current datasets, where sentences typically contain fewer than 10 relational triples and we’ve allocated sufficient slots, it could become problematic for more complex and longer documents. For scenarios with substantially more relation triples or other NLP objects (i.e., sets of mentions / POS / keywords), Hungarian matching becomes inefficient. Future implementations could explore Chamfer matching as an alternative, which offers reduced complexity while potentially maintaining adequate assignment quality.
@@ Appendix C Experiment Hyperparameter Configuration
Our experiments were conducted on a GeForce RTX 2080 Ti GPU (11GB VRAM) and ran with 12 random seeds (11, 29, 33, 39, 42, 53, 57, 62, 65, 73, 96, 98). Training time per epoch is approximately 7 minutes for NYT and 1 minute for WebNLG.
@@
Appendix D Prompting Strategy for Relational Triple Extraction
We prompt the model as a relation-triple extractor with the following instructions:
1.
Given a sentence, extract all relational triples present in it.
2.
Each triple must be formatted as a Python dictionary with exactly three keys: "subject", "relation", and "object".
3.
The "relation" field must take its value from a fixed inventory of 24 (NYT) / 218 (WebNLG)* predefined relation types, reflecting the target ontology for our task.
*Different from our data statistics because we included all the relations.
4.
The model is instructed to output a Python-style list of these dictionaries.
An example of the prompt template is as follows:
“‘ You are a relation-triple extractor. Given the following sentence, extract all relational triples. Each triple must be a dict with exactly three keys: ’subject’, ’relation’, ’object’. The ’relation’ value must be ONE of the following 24 options: ’industry’, ’place_lived’, ’founders’, ’advisors’, ’children’, ’neighborhood_of’, ’people’, ’profession’, ’place_of_birth’, ’ethnicity’, ’teams’, ’company’, ’administrative_divisions’, ’place_of_death’, ’geographic_distribution’, ’major_shareholders’, ’nationality’, ’place_founded’, ’capital’, ’country’, ’religion’, ’major_shareholder_of’, ’contains’, ’location’
Sentence: """[INPUT TEXT HERE]""" Output ONLY a Python list of dictionaries. For example: [{’subject’:’Bobby Fischer’, ’relation’:’nationality’, ’object’:’Iceland’}, {’subject’:’Iceland’, ’relation’:’capital’, ’object’:’Reykjavik’}] ”’
The model’s response is then post-processed by parsing the output using Python’s ‘ast.literal_eval‘ to safely convert the textual list of dictionaries into a Python object. If the parsing fails, typically due to formatting errors or hallucinations, the prediction is discarded and no triples are extracted for that sample.
In our benchmark evaluations, we did not encounter any parsing failures. However, preliminary experiments with smaller models (i.e., those with fewer than 7 billion parameters) often resulted in hallucinated or malformed outputs that could not be parsed. While these results were not formally included in this study, they highlight the importance of model scale and instruction-following capabilities in zero-shot information extraction tasks.
Appendix E Licensing and Terms of Use
E.1 Dataset
@@
E.2 Model and Code Release
We will release the code after acceptance under MIT license.
E.3 Software Dependencies and Implementation
@@
E.4 Ethical and Legal Considerations
@@
E.5 Information About Use Of AI Assistants